how do you find the value of k so that (k^2 , k) is on the line 2x-5y=3
thanks!

Answer by stanbon(75887)

how do you find the value of k so that
(k^2 , k) is on the line 2x-5y=3
----------------------
Substitute for x and y and solve for "k".
---
2k^2 - 5k - 3 = 0
----
(k-3)(2k+1) = 0
k = 3 or k = -1/2
-----
Ans: (9,3) or (1/4,-1/2)
===========================
Cheers,
Stan H.